# Quantum Coherence of Topologically Frustrated Spin Chains

### v1

@@ I Introduction
Since their discovery, several models have shown phenomenologies, often very different from each other, which can be encapsulated into the class of topological effects. Among them, the topologically frustrated (TF) models have attracted considerable interest in recent years [6, 7, 8]. This phenomenon was first identified in one-dimensional antiferromagnetic spin chains (AFMs) with nearest-neighbor coupling and frustrated boundary conditions (FBCs) [9, 10]. FBCs occur when periodic boundary conditions (PBCs) are imposed on a system with an odd number of spins. This configuration breaks the periodicity of the AFM interaction, introducing a delocalized excitation into the system. As a result of the presence of this excitation, the static [11, 12] and dynamical [13] properties of TF systems acquire distinctive and universal behaviors. These results aroused considerable interest, which further increased when it was shown that the phenomenon of topological frustration extends beyond the class of models in which it was initially unveiled. A notable example is the so-called ANNNI model [14], characterized by competing interactions between couples of spins at nearest and next-nearest neighbor sites along the lattice. In this model, topological frustration occurs when a system with an even number of spins, not divisible by four, is subjected to periodic boundary conditions, resulting in the emergence of two delocalized excitations.
@@
Among the various effects of imposing topological frustration on a one-dimensional system, one of the most surprising involves two fundamental quantum resources [15] such as the entanglement [16] and the non-stabilizerness (also known as magic) [17]. Specifically, in the frustrated phase [10, 18, 6, 19], as the chain length diverges, the values of these resources can be decomposed into two distinct contributions. The first is a somewhat local term, corresponding to the value the resource would assume when topological frustration is removed, i.e. when a different set of boundary conditions is applied to the system. In contrast, the second term has a topological nature and, therefore, remains independent of the specific Hamiltonian parameters. Thus, its value, which remains constant throughout the phase, can be evaluated (analytically) near the classical point, namely the point at which the Hamiltonian reduces to a sum of mutually commuting terms and the ground state can be calculated perturbatively. This contribution to the quantum resources can be attributed to the presence of delocalized excitations induced by TF, which are known to be responsible for the topological nature of the TF phase, as proved through the Disconnected Entanglement entropy (DEE) [20]. Indeed, a finite DDE, such as that found in [20], is a unique signature of the long-range correlation typical of topological phases, although a characteristic topological invariant has not been identified yet for TF.
@@
It is therefore natural to ask how general this picture is and whether or not it extends to other quantum resources. Unfortunately, various quantum resources can have different structures with few points of contact, and thus it does not seem feasible to provide a general proof of this property or to perform a broad analysis. For this reason, accepting our limits, in this paper we test the hypothesis of the generality of this picture by focusing on a further resource: Quantum Coherence (QC) [21, 22] The choice of the QC is not accidental but takes into account its relevance in several fields of research. Indeed, QC plays a fundamental role across multiple areas of quantum physics. In quantum information, QC is a key resource for quantum computation and communication, enabling quantum speedups and secure protocols [21, 23]. In quantum thermodynamics, QC influences work extraction, quantum fluctuation theorems, and the fundamental limits of energy transfer in nanoscale systems [24, 25, 26, 27]. In quantum many-body systems, its analysis can underlie the presence of quantum phase transitions, and the emergence of macroscopic quantum phenomena [28, 22, 29, 30]. Moreover, QC is essential for understanding the interplay between entanglement and non-classical correlations in open quantum systems [31].
To achieve the objective of this paper, it is necessary to disentangle the two contributions to QC. The topological contributions to QC are derived purely analytically by considering the system in proximity of the classical point. All other data are obtained through numerical computations. Given the exponential complexity of the evaluation, we employ the Density Matrix Renormalization Group (DMRG) algorithm to efficiently represent the ground state of the system and the tensor-cross-interpolation (TCI) algorithm to construct an efficient representation of QC as a univariate function [32]. Our results indicate that as the system size increases, the QC in the frustrated phase becomes increasingly well approximated by the sum of the topological and the local terms. This observation, together with previous findings for entanglement entropy and stabilizer Rényi entropy, suggests a universal behavior for at least all entropy-based quantum resource quantifiers.
@@
The paper is structured as follows. In Sec.II, we provide a brief overview of the numerical approach proposed in [32] for computing the QC as measured by the Relative Entropy of Coherence (REC). In Sec. III, we describe the models under consideration and summarize their known phenomenology. Section IV presents our findings, demonstrating that, similar to entanglement and magic, REC decomposes into two distinct contributions: a local term, corresponding to the REC value in the absence of TF, and a second contribution of topological origin. Finally, in Sec. V we summarize the results and discuss future developments, and in Appendix A we provide the analytical calculation of REC for the ground state of the ANNNI chain.
@@ II Measuring Quantum Coherence
Over the years, several different measures of QC have been suggested [22], each of them with its positive and negative sides. Since we are interested solely in pure states, in this work, we resort to the Relative Entropy of Coherence (REC) [21]. For a state with a density matrix $\rho$ the REC is defined as $C(\rho)=S(\rho_{\textrm{diag}})-S(\rho)$, where $S(\rho)$ stands for the von Neumann entropy and $\rho_{\textrm{diag}}$ is the matrix obtained by $\rho$ by deciding on a reference basis frame and setting all off-diagonal elements in this basis to zero. For a pure state, i.e. for $\rho=\ket{\psi}\!\bra{\psi}$ |, $S(\rho)$ vanishes and the REC reduces to
@@ III Models
Using REC, we tested our hypothesis on quantum resources on two different one-dimensional models that can, under suitable conditions, show the presence of TF. Both these models can be obtained for a particular choice of parameters, from the following Hamiltonian
@@
Setting $J_{2}=0$ we recover the well-known Ising model with a transverse field [41, 42] that shows a quantum phase transition at $|h|=h_{c}\equiv 1$. Choosing $J_{1}>0$ and imposing FBCs, for $0<|h|<h_{c}$, the system is in a topologically frustrated phase with a gapless energy spectrum in the thermodynamic limit. For $h=0$ (the classical point mentioned above), there are $2L$ exactly degenerate lowest energy states, namely the mutually orthogonal kink states |k+$\ket{k^{+}}$ and |$\ket{k^{-}}$, which are individually separable and diagonal in the computational basis. Indeed, because of PBCs and $L$ being odd, it is impossible to have perfect Neel (staggered) order these states present a magnetic defect, i.e. two parallel aligned neighboring spins, which can be placed on any bond of the ring. Hence |k+$\ket{k^{+}}$ and |$\ket{k^{-}}$ are the two states with the defect between the $k$-th and the $k+1$-th sites and differ from each other by the orientation of the defect. A finite magnetic field hybridizes these states and lifts the degeneracy, resulting in the aforementioned energy band. In particular, in the limit $h\rightarrow 0^{+}$ the GS is given by |$\ket{g}=\frac{1}{\sqrt{2L}}\sum_{k=1}^{L}(\ket{k^{+}}+\ket{k^{-}})$ and the whole band can be characterized as a single delocalized kink excitation with some momentum.
@@
On the contrary, considering $J_{2}>0$, we recover the so-called ANNNI model in which, independently of boundary conditions and the sign of $J_{1}$, there is always geometrical frustration induced by the competition between the two interactions. The phase diagram of this model is rich, having four distinct phases depending on the dominant interaction and the value of the magnetic field [14]. Among these, one of the most peculiar is the so-called antiphase which can occur when the next-to-near neighbor interaction $J_{2}$ dominates over $J_{1}$, i.e. if and only if $\kappa=|J_{1}|/J_{2}>1/2$. In such a phase, TF can be induced considering a ring made of an even number of spins that is not an integer multiple of four ($L=4n+2,n\in\mathbb{N}$). On the classical line $h=0$, the presence of TF increases the degeneration of the ground state manifold, which goes from a dimension equal to 4 (in the absence of TF) to $L^{2}/2$ if $J_{1}\neq 0$ or to $L^{2}$ in the limiting case $J_{1}=0$, where the chain decomposes into two independent subchains for the even and odd sites. Turning on a small transverse field, the degeneracy is lifted and the system admits a unique ground state given in Eq. (4), that can be described in terms of a distance-modulated superposition of kink states [19] (see Appendix A for details). Regardless of the presence or absence of the TF, by fixing the value of the ratio $\kappa$ and further increasing the value of $h$, the system undergoes two successive phase transitions. In the first, the system exits the antiphase and enters the so-called floating phase, whose nature has not yet been fully understood. It is characterized by a violation of the area law in terms of entanglement and is conjectured to be in a Luttinger liquid phase with algebraic correlations. Subsequently, the second transition marks the passage from the floating phase to a paramagnetic phase, where the dynamic is dominated by the transverse field.
@@
Figure 1: Upper Panel: Quantum Coherence of the ground state of the TF Ising chain (red full markers) and of the corresponding unfrustarted FM model (blue empty markers) as a function of the magnetic field and for different values of the system size. Middle Panel: The same quantities are plotted as a function of the system length for different values of the magnetic field. Bottom Panel: The ratio $R$ in eq. (3) for the Ising model as a function of the magnetic field and for different system sizes.
@@ IV Results
Having introduced the models of interest, let us discuss our results for the calculation of REC on the computational basis. It is crucial to emphasize that the choice of basis for calculating QC is fundamental, as QC, like magic but unlike entanglement, is a basis-dependent resource. Among various options, we choose to work on the computational basis because, near the classical point and in the absence of topological frustration, quantum coherence reaches its minimum.
@@
In Fig.1, we present our findings for the Ising model $(J_{2}=0)$. In the upper panel, we display the raw data obtained for the REC as a function of the applied magnetic field, both with $(J_{1}=1)$ and without $(J_{1}=-1)$ TF. Although our data are derived from systems of finite sizes, two distinct behaviors corresponding to the two phases of the model are observable. For large magnetic field $h$, the REC of the TF and non-TF Ising model converges to the same value. On the other hand, when the interaction term dominates over the external field, the behaviors differ significantly, and the difference increases with the system size. In particular, the limit $h\rightarrow 0^{+}$ is of special interest in this region. For the unfrustrated system, as $h$ approaches zero, $C(\rho)$ becomes independent of $L$ and tends to 1, consistent with the fact that the ground state can be well approximated by the GHZ state |$\ket{\phi}=\frac{1}{\sqrt{2}}(\ket{\uparrow}^{\otimes L}+\ket{\downarrow}^{\otimes L})$, where $\ket{\uparrow}$ and $\ket{\downarrow}$ are the eigenstates of $\sigma^{z}$ with eigenvalues $+1$ and $-1$, respectively. In contrast, with TF, for $h\to 0^{+}$ the REC tends to $1+\log_{2}(L)$, consistently with what can be inferred from the expression for the ground state of the TF Ising model presented in the previous section.
@@
Increasing $h$ but remaining in the same phase, the independence on the size disappears and the unfrustrated system shows a volume law, which recalls the behavior of magic, as measured by the stabilizerness Rényi entropy [6], as can be seen in the middle panel of Fig. 1. The analogy between the stabilizer Rényi entropy and the REC holds in TF systems as well, where both exhibit a volume-law behavior in the large-size limit, although with a significant correction for smaller systems.
To test our hypothesis that, in the thermodynamic limit, any quantum resource evaluated for the ground state of a TF model becomes the sum of a local term and a topological one that does not depend on the Hamiltonian parameters, we adopt the same approach as used in [18] for the Entanglement Entropy and in [6, 8] for the Stabilizer R’enyi Entropy. This approach involves the study of the ratio
$R=\frac{Q(\rho_{\textrm{fr}})-Q(\rho_{\textrm{unfr}})}{Q(\rho_{\textrm{fr}}^{h\rightarrow 0^{+}})-Q(\rho_{\textrm{unfr}}^{h\rightarrow 0^{+}})},$ (3)
where $Q$ represents an entropic measure of a generic quantum resource, while $\rho_{\textrm{fr}}$ and $\rho_{\textrm{unfr}}$ denote the density matrices of the ground states for the topologically frustrated and topologically unfrustrated models, respectively, given the same Hamiltonian parameters.
@@
If our initial hypothesis holds, the ratio in Eq. (3) should converge to 1 in the frustrated phase and approach zero in the paramagnetic phase. To compute this ratio, we first need to determine the denominator. For the Ising model, as discussed in the previous section, near the classical point and in the presence of topological frustration, the ground state can be approximated as a homogeneous superposition of $2L$ elements from the computational basis, yielding a REC of $\log_{2}(2L)$. Conversely, in the same region of parameter space but without frustration, the ground state simplifies to a GHZ state with a REC of 1. Thus, the denominator of Eq. (3) reduces to $\log_{2}(L)$. The bottom panel of Fig.1 shows our numerical results for $R$ in the case of the Ising model as a function of the magnetic field. Consistent with our working hypothesis, we observe that, as $L$ diverges, the value approaches $1$ throughout the whole topological phase and zero in the paramagnetic phase. The speed at which these values are reached depends on the ratio between the correlation length and the system size. Since the correlation length diverges at the phase transition, finite-size effects vanish more rapidly the further one moves away from $h=1$. We can then write the REC in the thermodynamic limit as the sum of the topological term and the magnetic one.
@@
Figure 2: Upper Panel: Quantum Coherence of the ground state of the TF ANNNI chain (red full markers) and of the corresponding unfrustarted OBCs model (blue empty markers) as a function of the magnetic field and for different values of the system size. Middle Panel: The same quantities are plotted as a function of the system length for different values of the magnetic field. Bottom Panel: The ratio $R$ in eq. (3) for the ANNNI model as a function of the magnetic field and for different system sizes.
@@
Let us now consider the ANNNI model by introducing an antiferromagnetic next-to-nearest-neighbor interaction ($J_{2}>0$). Even if the two models seem not so different from each other, the ANNNI model is much more complex to analyze, and several technical challenges must be addressed. The first problem arises when trying to associate the specific non-TF model with its TF equivalent. For the Ising model, this link is natural. Indeed, assuming periodic boundary conditions (PBCs) and an odd number of spins, transitioning between TF and non-TF Ising systems can be achieved simply by inverting the sign of $J_{1}$. This approach does not apply to ANNNI models. Specifically, when PBCs are imposed and the number of spins is even but not an integer multiple of 4, the presence of the TF is unaffected by the sign of $J_{1}$. On the other hand, by changing the sign of $J_{2}$, we not only remove TF but also the extensive frustration characteristic of the model, thereby fundamentally altering the properties of the system. Therefore, according to [19], this comparison is done by opening up the ring and moving from PBCs to open boundary conditions (OBCs). This introduces finite-size effects in the unfrustrated system which are expected to become irrelevant in the thermodynamic limit.
@@
The lower panel of Fig.2 shows the behavior of the ratio in (3) as a function of $h$ for a fixed value of $\kappa=|J_{1}|/J_{2}>1/2$ and allows us to test our working hypothesis. As we can see, despite the relatively small size of the system considered, constrained by the large bond size in the DMRG algorithm required to obtain an accurate estimate of the ground state, the thermodynamic behavior is evident. In the TF phase, the ratio Eq. (3) approaches 1 while, in the paramagnetic phase, it drops to 0 in a complete analogy with the Ising model.

### v2

@@ I Introduction
Since their discovery, several models have shown phenomenologies, often very different from each other, which can be encapsulated into the class of topological effects. Among them, the topologically frustrated (TF) models have attracted considerable interest in recent years [6, 7, 8, 9, 10, 11, 12, 13, 14]. This phenomenon was first identified in one-dimensional antiferromagnetic spin chains (AFMs) with nearest-neighbor coupling and frustrated boundary conditions (FBCs) [14, 15]. FBCs occur when periodic boundary conditions (PBCs) are imposed on a system with an odd number of spins. This configuration breaks the periodicity of the AFM interaction, introducing a delocalized excitation into the system. A similar phenomenology has also been identified in chains with even-number sites and tuning the frustration on and off by switching between periodic and anti-periodic boundary conditions [16]. As a result of the presence of this excitation, the static [17, 18] and dynamical [19] properties of TF systems acquire distinctive and universal behaviors that can be exploited for quantum technological applications [10, 8, 20]. Moreover, this phenomenology extends far beyond the original class of models in which it was first discovered. A prominent example is the ANNNI model [21], which features competing interactions between spins located at nearest and next-nearest neighbor sites along the lattice. In this system, topological frustration arises under periodic boundary conditions when the total number of spins is even but not divisible by four, leading to the appearance of two delocalized excitations [22].
@@
Among the various effects of imposing topological frustration on a one-dimensional system, one of the most surprising involves two fundamental quantum resources [23] such as the entanglement [24] and the non-stabilizerness (also known as magic) [25]. Specifically, in the frustrated phase [15, 26, 6], as the chain length diverges, the values of these resources can be decomposed into two distinct contributions. The first is a somewhat local term, and stays unchanged even if the topological frustration is removed, i.e., when a different set of boundary conditions is applied to the system. In contrast, the second term has a topological nature and, therefore, remains independent of the specific Hamiltonian parameters. Thus, its value, which remains constant throughout the phase, can be evaluated near the classical point, namely the point at which the Hamiltonian reduces to a sum of mutually commuting terms, and the ground state can be calculated perturbatively. This contribution to the quantum resources can be attributed to the presence of delocalized excitations induced by TF, which are known to be responsible for the topological nature of the TF phase, as proved through the Disconnected Entanglement entropy (DEE) [27]. A finite DEE is a unique signature of the long-range correlation typical of topological phases, although a characteristic topological invariant has not been identified yet for TF.
@@
It is therefore natural to ask how general this picture is and whether it extends to other quantum resources. Unfortunately, various quantum resources can have different structures with few points of contact, and thus it does not seem feasible to provide a general proof of this property or to perform a broad analysis. For this reason, in this paper we test the hypothesis of the generality of this picture by focusing on a further resource: Quantum Coherence (QC) [28, 29]. The choice of the QC is not accidental but takes into account its relevance in several fields of research. Indeed, QC plays a fundamental role across multiple areas of quantum physics. In quantum information, QC is a key resource for quantum computation and communication, enabling quantum speed-ups and secure protocols [28, 30]. In quantum thermodynamics, QC influences work extraction, quantum fluctuation theorems, and the fundamental limits of energy transfer in nanoscale systems [31, 32, 33, 34]. In quantum many-body systems, its analysis can underlie the presence of quantum phase transitions, and the emergence of macroscopic quantum phenomena [35, 29, 36, 37, 38]. Moreover, QC is essential for understanding the interplay between entanglement and non-classical correlations in open quantum systems [39].
This paper presents an analysis based on a DMRG-driven numerical approach, which employs tensor networks to efficiently represent the ground state of the system. To construct a compact and accurate representation of quantum coherence (QC) as a univariate function, we also make use of the Tensor Cross Interpolation (TCI) algorithm [40]. As is well established, QC is a basis-dependent quantum resource. Since it is unfeasible to analyze all possible bases in the Hilbert space of a large qubit system, our investigation is primarily conducted in the computational basis. This choice is motivated by two main reasons. First: near the classical point and in absence of topological frustration, the ground state approaches a GHZ-like form in this basis, where quantum coherence reaches a minimum. Second: prior studies on quantum magic in these models have also used the computational basis, allowing for a more direct comparison between the behavior of coherence and magic. The results of our analysis are consistent with earlier findings on entanglement entropy and non-stabilizerness entropy, suggesting the possibility of a universal structure shared by entropy-based quantum resource measures.
@@
The paper is structured as follows. In Sec.II, we provide a brief overview of the numerical approach proposed in [40] for computing the QC as measured by the Relative Entropy of Coherence (REC). In Sec. III, we describe the models under consideration and summarize their known phenomenology. Section IV presents our results, demonstrating that, similar to entanglement and magic, for large systems, the REC in the computational basis decomposes into two distinct contributions: one is local, and the other is of topological origin. Finally, in Sec. V we summarize the results and discuss future developments. Additional important results are presented in the appendices. Appendix A contains the analytical computation of the REC in the computational basis for the ground state of the ANNNI chain near the classical point. In Appendix B, we present our numerical analysis considering other bases generated by rigid rotations and show that our findings, that is the decomposition of REC between a local and a topological contribution, are valid beyond the computational basis on which we concentrated in the bulk of the manuscript. Lastly, Appendix C addresses the numerical challenges associated with these evaluations.
@@ II Measuring Quantum Coherence
Over the years, several different measures of QC have been suggested [29], each of them with its positive and negative sides. Since we are interested solely in pure states, in this work, we resort to the Relative Entropy of Coherence (REC) [28]. For a state with a density matrix $\rho$ the REC is defined as $C(\rho)=S(\rho_{\textrm{diag}})-S(\rho)$, where $S(\rho)$ stands for the von Neumann entropy and $\rho_{\textrm{diag}}$ is the matrix obtained by $\rho$ by deciding on a reference basis frame and setting all off-diagonal elements in this basis to zero. For a pure state, i.e., for $\rho=\ket{\psi}\!\bra{\psi}$ |, $S(\rho)$ vanishes and the REC reduces to
@@ III Models
Using REC, we tested our hypothesis on quantum resources of two different one-dimensional models that can, under suitable conditions, show the presence of TF. Both these models can be obtained for a particular choice of parameters, from the following Hamiltonian
@@
Setting $J_{2}=0$ we recover the well-known Ising model with a transverse field [49, 50] that shows a quantum phase transition at $|h|=h_{c}\equiv 1$. Choosing $J_{1}>0$ and imposing FBCs, for $0<|h|<h_{c}$, the system is in a topologically frustrated phase with a gapless energy spectrum in the thermodynamic limit. For $h=0$ (the classical point mentioned above), there are $2L$ exactly degenerate lowest energy states, namely the mutually orthogonal kink states |k+$\ket{k^{+}}$ and |$\ket{k^{-}}$, which are individually separable and incoherent in the computational basis. Indeed, because of PBCs and $L$ being odd, it is impossible to have perfect Neel (staggered) order and therefore these states present a magnetic defect, i.e. two parallel aligned neighboring spins, which can be placed on any bond of the ring. Hence |k+$\ket{k^{+}}$ and |$\ket{k^{-}}$ are the two states with the defect between the $k$-th and the $k+1$-th sites and differ from each other by the orientation of the defect. Switching on a finite magnetic field (i.e., setting $h>0$) leads to a hybridization of these states, thereby lifting their degeneracy and giving rise to the resulting (gapless) energy band. In particular, in the limit $h\rightarrow 0^{+}$ the GS is given by |$\ket{g}=\frac{1}{\sqrt{2L}}\sum_{k=1}^{L}(\ket{k^{+}}+\ket{k^{-}})$ and the whole band can be characterized as a single delocalized kink excitation with some momentum.
@@
On the contrary, considering $J_{2}>0$, we recover the so-called ANNNI model in which, independently of boundary conditions and the sign of $J_{1}$, there is always geometrical frustration induced by the competition between the two interactions. The phase diagram of this model is rich, featuring four distinct phases that depend on the dominant interaction and the value of the magnetic field [21]. Among these, one of the most peculiar is the so-called antiphase which can occur when $\kappa=|J_{1}|/J_{2}>1/2$. In such a phase, TF can be induced considering a ring made of an even number of spins that is not an integer multiple of four ($L=4n+2,n\in\mathbb{N}$). On the classical line $h=0$, the presence of TF increases the degeneracy of the ground state manifold, which goes from a dimension equal to 4 (in the absence of TF) to $L^{2}/2$ if $J_{1}\neq 0$ or to $L^{2}$ in the limiting case $J_{1}=0$, where the chain decomposes into two independent subchains for the even and odd sites. Turning on a small transverse field, the degeneracy is lifted and the system admits a unique ground state given in Eq. (4), that can be described in terms of a distance-modulated superposition of kink states [22] (see Appendix A for details). Regardless of the presence or absence of the TF, by fixing the value of the ratio $\kappa$ and further increasing the value of $h$, the system undergoes two successive phase transitions. In the first, the system exits the antiphase and enters the so-called floating phase, whose nature has not yet been fully understood. It is characterized by a violation of the area law in terms of entanglement and is conjectured to be in a Luttinger liquid phase with algebraic correlations. Subsequently, the second transition marks the passage from the floating phase to a paramagnetic phase, where the dynamic is dominated by the transverse field.
@@
Figure 1: (a) Quantum Coherence of the ground state of the TF Ising chain (red full markers) and of the corresponding unfrustrated FM model (blue empty markers) as a function of the magnetic field and for different values of the system size. (b) The same quantities are plotted as a function of the system length for different values of the magnetic field. (c) The ratio $R$ in eq. (3) for the Ising model as a function of the magnetic field and for different system sizes.
@@ IV Results
Having introduced the models under consideration and the TCI-based method for computing the REC, we now proceed to present and discuss our results. In this section, we focus specifically on evaluating the REC in the computational basis, while a generalization to a broader class of bases is provided in Appendix B. The computational basis comprises states that are tensor products of the eigenstates of $\sigma^{z}_{i}$, where $i$ labels the sites of the system. Accordingly, each basis state can be expressed as |$\ket{\phi}=\bigotimes_{i=1}^{N}\ket{\phi_{i}}$, with each |$\ket{\phi_{i}}$ equal to either |$\ket{\uparrow_{i}}$ or |$\ket{\downarrow_{i}}$. Once the RECs have been evaluated, we test our hypothesis that, in the thermodynamic limit, any quantum resource computed for the ground state of a TF model decomposes into a sum of a local term and a topological term, the latter independent of the specific Hamiltonian parameters. To this end, we employ the same approach previously used in [26] for Entanglement Entropy and in [6, 7] for the Stabilizer Rényi Entropy. This method involves analyzing the ratio
$R=\frac{C(\rho_{\textrm{fr}})-C(\rho_{\textrm{unfr}})}{C(\rho_{\textrm{fr}}^{h\rightarrow 0^{+}})-C(\rho_{\textrm{unfr}}^{h\rightarrow 0^{+}})},$ (3)
where $\rho_{\textrm{fr}}$ and $\rho_{\textrm{unfr}}$ denote the ground state density matrices of the topologically frustrated and unfrustrated models, respectively, characterized by having the same correlation length. According to our hypothesis, these two states share an identical local contribution to the QC in the thermodynamic limit, while differing in their topological terms. Thus, subtracting one from the other isolates the difference between the topological contributions. Similarly, $\rho_{\textrm{fr}}^{h\rightarrow 0^{+}}$ and $\rho_{\textrm{unfr}}^{h\rightarrow 0^{+}}$ correspond to the respective ground state density matrices near the classical limit, as $h\rightarrow 0^{+}$, where the Hamiltonian comprises mutually commuting terms. If our hypothesis holds, the difference in coherence between these two states should likewise capture just the difference between the topological components. As a result, the ratio defined in Eq. (3) compares two independent estimates of the same topological difference. If this difference is indeed topological in nature, the ratio should remain invariant under variations of the system parameters, provided the system stays within the same macroscopic phase. Therefore, if our hypothesis is valid—namely, that in the thermodynamic limit the QC of a topologically frustrated system splits into local and topological contributions—then the ratio in Eq. (3) should converge to 1 throughout the topologically ordered phase.
@@
In Fig.1, we present our findings for the Ising model $(J_{2}=0)$. In this case, to guarantee that the topologically frustrated and non-frustrated models share the same correlation length, it is sufficient to consider systems with periodic boundary conditions (PBC), composed of the same odd number of spins, subject to the same value of $h$, but with opposite values of $J$ — specifically, $+1$ for the frustrated model and $-1$ for the non-frustrated one. In the upper panel, we display the raw data obtained for the REC as a function of the applied magnetic field, both with and without TF. Although our data are derived from systems of finite sizes, two distinct behaviors corresponding to the two phases of the model are observable. For large magnetic fields $h$, the REC of the TF and non-TF Ising model converges to the same value. On the other hand, when the interaction term dominates over the external field, the behaviors differ significantly, and the difference increases with the system size. In particular, the limit $h\rightarrow 0^{+}$ is of special interest in this region. For the unfrustrated system, as $h$ approaches zero, $C(\rho)$ becomes independent of $L$ and tends to 1, consistent with the fact that the ground state can be well approximated by the GHZ state |$\ket{\phi}=\frac{1}{\sqrt{2}}(\ket{\uparrow}^{\otimes L}+\ket{\downarrow}^{\otimes L})$, where $\ket{\uparrow}$ and $\ket{\downarrow}$ are the eigenstates of $\sigma^{z}$ with eigenvalues $+1$ and $-1$, respectively. In contrast, with TF, for $h\to 0^{+}$ the REC tends to $1+\log_{2}(L)$, consistently with what can be inferred from the expression for the ground state of the TF Ising model presented in the previous section.
@@
Increasing $h$ but remaining in the same phase, the independence on the size disappears and the QC in the unfrustrated system shows a volume-law, which recalls the behavior of magic, as measured by the stabilizerness Rényi entropy [6], as can be seen in the middle panel of Fig. 1. The analogy between the stabilizer Rényi entropy and the REC holds in TF systems as well, where both exhibit a volume-law behavior in the large-size limit, although with a significant correction for smaller systems.
@@
The bottom panel of Fig.1 shows our numerical results for $R$ in the case of the Ising model as a function of the magnetic field. Consistent with our working hypothesis, we observe that, as $L$ diverges, the value approaches $1$ throughout the whole topological phase and zero in the paramagnetic phase. The speed at which these values are reached depends on the ratio between the correlation length and the system size. Since the correlation length diverges at the phase transition, finite-size effects vanish more rapidly the further one moves away from $h=1$. We can then write the REC in the thermodynamic limit as the sum of the topological term and the magnetic one.
@@
Figure 2: (a) Quantum Coherence of the ground state of the TF ANNNI chain (red full markers) and of the corresponding unfrustrated OBCs model (blue empty markers) as a function of the magnetic field and for different values of the system size. (b) The same quantities are plotted as a function of the system length for different values of the magnetic field. (c) The ratio $R$ in eq. (3) for the ANNNI model as a function of the magnetic field and for different system sizes.
@@
Let us now consider the ANNNI model by introducing an antiferromagnetic next-to-nearest-neighbor interaction ($J_{2}>0$). Even if the two models seem not so different from each other, the ANNNI model is much more complex to analyze, and several technical challenges must be addressed. The first problem arises when trying to associate the specific non-TF model with its TF equivalent, i.e., when we try to select the two states (one frustrated and one unfrustrated) that share the same correlation length. For the Ising model, this link is natural. Indeed, assuming periodic boundary conditions (PBCs) and an odd number of spins, transitioning between TF and non-TF Ising systems can be achieved simply by inverting the sign of $J_{1}$. This approach does not apply to ANNNI models. Specifically, when PBCs are imposed and the number of spins is even but not an integer multiple of 4, the presence of the TF is unaffected by the sign of $J_{1}$. On the other hand, by changing the sign of $J_{2}$, we not only remove TF but also the extensive frustration characteristic of the model, thereby fundamentally altering the properties of the system. Therefore, according to [22], this comparison is done by opening up the ring and moving from PBCs to open boundary conditions (OBCs). This introduces finite-size effects in the unfrustrated system, which are expected to become irrelevant in the thermodynamic limit.
@@
The bottom panel of Fig.2 shows the behavior of the ratio in (3) as a function of $h$ for a fixed value of $\kappa=|J_{1}|/J_{2}>1/2$ and allows us to test our working hypothesis. As we can see, despite the relatively small size of the system considered, constrained by the large bond size in the DMRG algorithm required to obtain an accurate estimate of the ground state, the thermodynamic behavior is evident. In the TF phase, the ratio Eq. (3) approaches 1 while, in the paramagnetic phase, it drops to 0 in a complete analogy with the Ising model.